OpenStudy (anonymous):

If a circle has a circumference of 14 cm, then what is the radius of the circle? A. 7 cm B. 7 cm C. 2 cm D. 14 cm

OpenStudy (anonymous):

circumference = 2*pi*r^2

OpenStudy (anonymous):

solve for r by dividing by 2pi first... So 14/2pi = 2.228

OpenStudy (anonymous):

Now we know r^2=2.228 then you need to take the square root of 2.228 which is around 1.493

OpenStudy (anonymous):

And that's your radius

OpenStudy (anonymous):

would it be horrible to say I dont understand? ;~;

OpenStudy (anonymous):

circum =2pir .. 14=2pir ... divide both sides by 2pi ..